OpenStudy (anonymous):

a 150g ball moving at 0.45 m/s in the +x direction is subject to a 0.15N force in the -y direction for 1.5s. find the ball's momentum change and velocity

OpenStudy (anonymous):

to find momentum i use this equation\[\sum_ 1 F*\Delta t=\Delta p\] and the Force is negative since it is going in a negative direction so it will be -0.15kg * 1.5= -0.225 kgm/s since i found the momentum change i use this equation to find velocity \[\Delta p=m(v-v_0)\] -0.225=0.15(v-0.45) v=-1.05m/s

OpenStudy (anonymous):

the answers were incorrect and i do not know what i am doing wrong.

OpenStudy (jamesj):

Remember that this force is perpendicular to the direction of motion of the object and hence gives it a new component of momentum. Before the force interacts, the momentum vector is p = mv = 0.15(0.45, 0) = (0.00675,0) The impulse of the force is Ft = (0.15)(1.5) = 0.225 in the j/y direction, hence the momentum vector afterwards is p = (0.00675, 0.225)

OpenStudy (jamesj):

*negative y direction hence the new momentum is p = (0.00675, -0.225)

OpenStudy (jamesj):

Now back out mass again to find the velocity vector v, p = mv

OpenStudy (anonymous):

what do you mean by "back out"

OpenStudy (jamesj):

I mean that to find the velocity vector, v, you will need to divide the component of the momentum vector, p, by the scalar mass, m.

OpenStudy (jamesj):

that is, if momentum is p = (p1, p2), then the velocity is v = (p1/m, p2/m) = (v1, v2)

OpenStudy (anonymous):

okay. i get it thank you.